OpenStudy (anonymous):

calculate the value of k such that kx^2-4x=0 has one root. Steps if possible, thank you very much

OpenStudy (amistre64):

b^2-4ac = 0 has only one root

OpenStudy (amistre64):

well, its called a repeated root, but its only one value

OpenStudy (amistre64):

divide by k; so that you have x^2 -(4/k)x = 0 x(x-(4/k))= 0 when x=0 and x=4/k i dont think there you will find a value such that you get only one root unless you make it a linear

OpenStudy (anonymous):

thank you very much